OpenStudy (anonymous):

Calculate the number of molecules present in 3.50 mol FeO2. 2.11 x 1024 molecules 2.107 x 1023 molecules 3.50 x 1023 molecules

Vocaloid (vocaloid):

1 mole = 6.022*10^23 molecules 3.50 mole = 3.50*6.022*10^23 molecules

OpenStudy (anonymous):

thank u